The molar mass of water is 18.0 grams per mole. The % of oxygen in water is:


1
Expert's answer
2020-12-21T03:48:09-0500

Molar mass of oxygen = 16

Molar mass of water = 18

So percentage = 16/18 = 0.89(100) = 89%